# Green’s-Function Spherical Neural Operators for Biological Heterogeneity

### v2

@@ Abstract
Spherical deep learning has been widely applied to a broad range of real-world problems. Existing approaches often face challenges in balancing strong spherical geometric inductive biases with the need to model real-world heterogeneity. To solve this while retaining spherical geometry, we first introduce a designable Green’s function framework (DGF) to provide new spherical operator solution strategy: Design systematic Green’s functions under rotational group. Based on DGF, to model biological heterogeneity, we propose Green’s-Function Spherical Neural Operator (GSNO) fusing 3 operator solutions: (1) Equivariant Solution derived from Equivariant Green’s Function for symmetry-consistent modeling; (2) Invariant Solution derived from Invariant Green’s Function to eliminate nuisance heterogeneity, e.g., consistent background field; (3) Anisotropic Solution derived from Anisotropic Green’s Function to model anisotropic systems, especially fibers with preferred direction. Therefore, the resulting model, GSNO can adapt to real-world heterogeneous systems with nuisance variability and anisotropy while retaining spectral efficiency. Evaluations on spherical MNIST, Shallow Water Equation, diffusion MRI fiber prediction, cortical parcellation and molecule structure modeling demonstrate the superiority of GSNO.
@@ 1 Introduction
This work aims to address the adaptability required to represent the complex, heterogeneous real-world spherical systems that do not adhere to the symmetry SO(3)-group. We propose the Green’s function Spherical Neural Operators (GSNO), fusing Equivariant, Invariant and Anisotropic components to simulate heterogeneous biological systems. We first establish a Designable Green’s Function Framework (DGF) to construct a new class of spherical operators. Based on this framework, we design three Green’s functions and derive the corresponding solution operator for different systems: (1) Equivariant Green’s function $G_{E}(R^{-1}u)$ based on the relative position $\rightarrow$ Existing Equivariant Operator for isotropic symmetry modeling; (2) Invariant Green’s function $G_{I}(u)$ based on the absolute position $\rightarrow$ A new Invariant Operator for invariance modeling; (3) Anisotropic Green’s function $G_{A}(u\cdot Rd)$ based on the learnable directions $\rightarrow$ A new Anisotropic Operator with learnable symmetry groups $SO(\theta)$ for anisotropy modeling. The three solution operators are concatenated to form GSNO for efficient heterogeneous modeling.
@@
    Three spherical operators derived from 3 Green’s functions, especially the new Invariant and Anisotropic Operators for biological heterogeneity.
@@ 2.1 Study design
As shown in the Figure 1 (C and D), we design Green’s-Function Spherical Neural Operator (GSNO), a new spectral learning fusing Equivariant, Invariant and Anisotropic Operators by concatenation to model biologically heterogeneous systems while retaining sphere-native.
@@ 2.1.1 Experimental design
We compare GSNO to other state-of-the-art methods for 5 spherical experimental scenarios with distinct system characteristics, involving various spherical sampling methods, resolutions, and application fields, to comprehensively evaluate the generalization of GSNO. For example, brain nerve fiber modeling and cortical parcellation emphasize GSNO’s ability to model biological heterogeneity, including invariance and anisotropy. And the different time scales and spatial scales in shallow water equation highlight the nonlinear and asymmetric modeling capability and stability of GSNO. Spherical digital recognition and atomic structure analysis demonstrate the generalization ability of GSNO for spherical symmetry systems. All experiments are implemented through PyTorch on 32GB V100 GPUs.
@@ 2.3.2 Cortical parcellation
Cortical parcellation requires both equivariance to handle large-scale spatial variations among individuals and biological heterogeneity modeling to capture nuisance variability and the local subtle features that determine the boundaries. This study evaluate the GSNO’s robustness to anatomical variability and its parcellation accuracy on the public Mindboggle-101 [klein2012101] and NAMIC-39 [desikan2006automated] datasets. The datasets comprises cortical surfaces with expert manual annotations, from which we extracted geometric features and mapped them onto the sphere following a preprocessing pipeline consistent with SPHARM-Net [desikan2006automated, ha2022spharm].
@@ 4.2.2 Invariant Operator from $G(u)$
Motivation: Spherical biomedical systems—such as cortical surface measurements, diffusion-derived features, or functional activation maps—are often subject to substantial variability arising from acquisition protocols, coordinate choices, and inter-subject alignment procedures. These sources of variability induce differences in the observed data that do not reflect underlying biological differences. For instance, global rotations of the sphere, scanner-dependent orientation conventions, or arbitrary surface parameterizations can significantly alter the numerical representation of a sample while leaving its biological interpretation unchanged. If left unconstrained, learning-based models may allocate substantial representational capacity to capturing such variability, thereby conflating coordinate-dependent artefacts with meaningful biological signals. This leads to reduced generalization, poor reproducibility across datasets, and sensitivity to irrelevant transformations. From a scientific standpoint, these effects are undesirable, as they obscure the relationship between the learned representation and the biological system of interest.
@@
Rather than attempting to learn invariance from data, the invariant operators impose it by construction. This guarantees that nuisance variability is systematically removed from the learned features, preventing the model from encoding information that is irrelevant to the task. As a result, representational capacity is redirected toward modelling biologically meaningful variation, improving robustness, interpretability, and cross-dataset generalization. In the context of spherical biomedical data, invariant operators thus serve as a principled mechanism for disentangling biological signal from acquisition- and coordinate-induced artefacts, ensuring that other learning is driven by other scientifically relevant information.
@@ 4.2.3 Anisotropic Operator from $G(u\cdot R\mathbf{d})$
Motivation: Neural tissue, cortical folding patterns, and white-matter fiber architectures exhibit pronounced directional structure that plays a fundamental role in their function and organization. These anisotropic systems with preferred directions exhibit different properties along different directions. Such anisotropy reflects direct heterogeneity of underlying biological mechanisms. Modelling these systems with operators that assume isotropic SO(3)-symmetry can impose overly restrictive symmetry assumptions, forcing the model to treat all rotations as equally valid transformations. In many real-world systems, however, directional preferences, local reference frames, or constrained rotational freedoms are intrinsic to the underlying physical or semantic structure. Enforcing full equivariance in such cases can suppress direction-dependent responses, reduce representational flexibility, and lead to the loss of critical anisotropic information. Anisotropic operators are motivated by the need to model systems whose symmetries are inherently partial, directional, or locally constrained, rather than globally isotropic while retaining sphere-native. They provide a controlled and interpretable way to incorporate biologically meaningful anisotropy into the model.
@@
where $\mathbf{d_{\theta}}$ is the learnable orientations. The anisotropic operators respect underlying geometric structure and allow the model to express distinct modes along distinct directions, reflecting the organization of biological substrates such as fiber tracts or cortical boundaries. Therefore, by explicitly modelling anisotropy where it is biologically warranted, the operators enable the capture of heterogeneous, structure-driven variation. They thus complement spherical operators by restoring expressivity in a principled and biologically grounded manner.

### v3

# Neural Operators for Biomedical Spherical Heterogeneity

@@ Abstract
Spherical deep learning has been widely applied to a broad range of real-world problems. Existing approaches often face challenges in balancing strong spherical geometric inductive biases with the need to model real-world heterogeneity. To solve this while retaining spherical geometry, we first introduce a designable Green’s function framework (DGF) to provide new spherical operator solution strategy: Design systematic Green’s functions under rotational group. Based on DGF, to model biomedical heterogeneity, we propose Green’s-Function Spherical Neural Operator (GSNO) fusing 3 operator solutions: (1) Equivariant Solution derived from Equivariant Green’s Function for symmetry-consistent modeling; (2) Invariant Solution derived from Invariant Green’s Function to eliminate nuisance heterogeneity, e.g., consistent background field; (3) Anisotropic Solution derived from Anisotropic Green’s Function to model anisotropic systems, especially fibers with preferred direction. Therefore, the resulting model, GSNO can adapt to real-world heterogeneous systems with nuisance variability and anisotropy while retaining spectral efficiency. Evaluations on spherical MNIST, Shallow Water Equation, diffusion MRI fiber prediction, cortical parcellation and molecule structure modeling demonstrate the superiority of GSNO.
@@ 1 Introduction
This work aims to address the adaptability required to represent the complex, heterogeneous real-world spherical systems that do not adhere to the symmetry SO(3)-group. We propose the Green’s function Spherical Neural Operators (GSNO), fusing Equivariant, Invariant and Anisotropic components to simulate heterogeneous biomedical systems. We first establish a Designable Green’s Function Framework (DGF) to construct a new class of spherical operators. Based on this framework, we design three Green’s functions and derive the corresponding solution operator for different systems: (1) Equivariant Green’s function $G_{E}(R^{-1}u)$ based on the relative position $\rightarrow$ Existing Equivariant Operator for isotropic symmetry modeling; (2) Invariant Green’s function $G_{I}(u)$ based on the absolute position $\rightarrow$ A new Invariant Operator for invariance modeling; (3) Anisotropic Green’s function $G_{A}(u\cdot Rd)$ based on the learnable directions $\rightarrow$ A new Anisotropic Operator with learnable symmetry groups $SO(\theta)$ for anisotropy modeling. The three solution operators are concatenated to form GSNO for efficient heterogeneous modeling.
@@
    Three spherical operators derived from 3 Green’s functions, especially the new Invariant and Anisotropic Operators for biomedical heterogeneity.
@@ 2.1 Study design
As shown in the Figure 1 (C and D), we design Green’s-Function Spherical Neural Operator (GSNO), a new spectral learning fusing Equivariant, Invariant and Anisotropic Operators by concatenation to model biomedically heterogeneous systems while retaining spherical geometry.
@@ 2.1.1 Experimental design
We compare GSNO to other state-of-the-art methods for 5 spherical experimental scenarios with distinct system characteristics, involving various spherical sampling methods, resolutions, and application fields, to comprehensively evaluate the generalization of GSNO. For example, brain nerve fiber modeling and cortical parcellation emphasize GSNO’s ability to model biomedical heterogeneity, including invariance and anisotropy. And the different time scales and spatial scales in shallow water equation highlight the nonlinear and asymmetric modeling capability and stability of GSNO. Spherical digital recognition and atomic structure analysis demonstrate the generalization ability of GSNO for spherical symmetry systems. All experiments are implemented through PyTorch on 32GB V100 GPUs.
@@ 2.3.2 Cortical parcellation
Cortical parcellation requires both equivariance to handle large-scale spatial variations among individuals and biomedical heterogeneity modeling to capture nuisance variability and the local subtle features that determine the boundaries. This study evaluate the GSNO’s robustness to anatomical variability and its parcellation accuracy on the public Mindboggle-101 [klein2012101] and NAMIC-39 [desikan2006automated] datasets. The datasets comprises cortical surfaces with expert manual annotations, from which we extracted geometric features and mapped them onto the sphere following a preprocessing pipeline consistent with SPHARM-Net [desikan2006automated, ha2022spharm].
@@ 4.2.2 Invariant Operator from $G(u)$
Motivation: Spherical biomedical systems—such as cortical surface measurements, diffusion-derived features, or functional activation maps—are often subject to substantial variability arising from acquisition protocols, coordinate choices, and inter-subject alignment procedures. These sources of variability induce differences in the observed data that do not reflect underlying biomedical differences. For instance, global rotations of the sphere, scanner-dependent orientation conventions, or arbitrary surface parameterizations can significantly alter the numerical representation of a sample while leaving its biomedical interpretation unchanged. If left unconstrained, learning-based models may allocate substantial representational capacity to capturing such variability, thereby conflating coordinate-dependent artefacts with meaningful biomedical signals. This leads to reduced generalization, poor reproducibility across datasets, and sensitivity to irrelevant transformations. From a scientific standpoint, these effects are undesirable, as they obscure the relationship between the learned representation and the biomedical system of interest.
@@
Rather than attempting to learn invariance from data, the invariant operators impose it by construction. This guarantees that nuisance variability is systematically removed from the learned features, preventing the model from encoding information that is irrelevant to the task. As a result, representational capacity is redirected toward modelling biomedically meaningful variation, improving robustness, interpretability, and cross-dataset generalization. In the context of spherical biomedical data, invariant operators thus serve as a principled mechanism for disentangling biomedical signal from acquisition- and coordinate-induced artefacts, ensuring that other learning is driven by other scientifically relevant information.
@@ 4.2.3 Anisotropic Operator from $G(u\cdot R\mathbf{d})$
Motivation: Neural tissue, cortical folding patterns, and white-matter fiber architectures exhibit pronounced directional structure that plays a fundamental role in their function and organization. These anisotropic systems with preferred directions exhibit different properties along different directions. Such anisotropy reflects direct heterogeneity of underlying biomedical mechanisms. Modelling these systems with operators that assume isotropic SO(3)-symmetry can impose overly restrictive symmetry assumptions, forcing the model to treat all rotations as equally valid transformations. In many real-world systems, however, directional preferences, local reference frames, or constrained rotational freedoms are intrinsic to the underlying physical or semantic structure. Enforcing full equivariance in such cases can suppress direction-dependent responses, reduce representational flexibility, and lead to the loss of critical anisotropic information. Anisotropic operators are motivated by the need to model systems whose symmetries are inherently partial, directional, or locally constrained, rather than globally isotropic while retaining sphere-native. They provide a controlled and interpretable way to incorporate biomedically meaningful anisotropy into the model.
@@
where $\mathbf{d_{\theta}}$ is the learnable orientations. The anisotropic operators respect underlying geometric structure and allow the model to express distinct modes along distinct directions, reflecting the organization of biomedical substrates such as fiber tracts or cortical boundaries. Therefore, by explicitly modelling anisotropy where it is biomedically warranted, the operators enable the capture of heterogeneous, structure-driven variation. They thus complement spherical operators by restoring expressivity in a principled and biomedically grounded manner.